OpenStudy (anonymous):

The sum of two consecutive integers is at least 185. What is the smaller of the two integers?

OpenStudy (imstuck):

If you have the sum of 2 consecutive integers, "sum" means you are adding, and consecutive means they come one after the other in counting. So let's call the first unknown number x, and the one right after it is x+1. Get that so far?

OpenStudy (gorv):

first integer =x next=x+1 x+x+1=185 2x+1=185

OpenStudy (anonymous):

mhmm

OpenStudy (imstuck):

Actually, @gorv, you are not using an equals sign here.

OpenStudy (anonymous):

well what is it then? i kinda need to know quickly. the right answer too.. plzz

OpenStudy (imstuck):

ok, then if we have x and x+1 as the integers and we know we have to add them, we're good so far. What we have to decide is what to do with that sum. The problem tells us that the sum is AT LEAST 185. So it's an inequality problem. Like this: